OpenStudy (anonymous):

What is the domain of f(x)=-3/x-4+2

OpenStudy (campbell_st):

so if the question \[f(x) = \frac{-3}{x - 4} + 2\]

OpenStudy (anonymous):

\[f(x)=-\frac{ 3 }{ x-4 }+2\]

OpenStudy (campbell_st):

ok... so the domain are what values can you input... look at the denominator of the fraction, what value of x gives a zero denominator? then the domain is all real x except that value

OpenStudy (michele_laino):

please note that you can not divide by zero

OpenStudy (michele_laino):

when for what value of x, we have x-4=0?